Rank the following difference scores and find the Wilcoxon T for this sample.

Difference Scores -8 2 -10 -4 7 -13 -17 -9
Ranks ? ? ? ? ? ? ? ?

T=?
Answer #1
difference scores Absolute difference Ranks signed ranks
-8 8 4 -4
2 2 1 1
-10 10 6 -6
-4 4 2 -2
7 7 3 3
-13 13 7 -7
-17 17 8 -8
-9 9 5 -5

The next step is to calculate the W+ and W-.

W+ = 1+ 3 =4

W- = 4 + 6 + 2 + 7 + 8 + 5 = -32

T = max (W+ ,W- ) = max ( 4, -32 ) = 4

T= 4